OpenStudy (anonymous):

Note: You may use the following formulas to help you complete the data table: °C = (°F - 32) 5/9 ; °F = °C 9/5 + 32; °K = °C + 273; °K = 5/9 (°F - 32) + 273 Planet Name Average Temperature (°F) Average Temperature (°C) Average Temperature (°K) Mercury Venus Earth Mars Jupiter Saturn Uranus Neptune please someone help me im helping my boy!

OpenStudy (anonymous):

I will give you medal!

OpenStudy (anonymous):

Hi boys and girls Im Mrs.Viviana and i need help with something im helping my boy but i dont understan can someone please help i will give medal ( :

OpenStudy (anonymous):

Is this question complete? Is that all the information that was given or does it exepct you to do research on the temperature of the planets? The formulas given help you convert a temperature, but you haven't given us any numbers.

OpenStudy (anonymous):

Okay well this is all they gave me A) Scientists have narrowed down the origin of the sample to one of the planets listed in the table below. Using your knowledge of temperature scales, find the appropriate Fahrenheit, Celsius, and Kelvin temperatures to complete the data table below. (12 points; ½ point per item) Note: You may use the following formulas to help you complete the data table: °C = (°F - 32) 5/9 ; °F = °C 9/5 + 32; °K = °C + 273; °K = 5/9 (°F - 32) + 273 Planet Name Average Temperature (°F) Average Temperature (°C) Average Temperature (°K) Mercury Venus Earth Mars Jupiter Saturn Uranus Neptune

OpenStudy (anonymous):

The question you wrote proves there's more. What does the question say about the "sample" or the other planets. If this is question A, then it must be under more information To solve the question, the question is asking to prove that the sample is from one of the planets. If that's all that's really said, then the only thing I can think of would be to use google and search for the temperature of each planet and use the formulas to convert them.

OpenStudy (anonymous):

ie from Google I found that the avg temp of mercury is 340 K Now use the formulas to convert K to C and C to F

OpenStudy (anonymous):

thank you very much now is really clear to me!